# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

### v1

@@ 4.1 Regularity analysis
In this section, we will investigate how to obtain a remainder term with Sobolev regularity $\frac{1}{2}+$ under weaker dissipation conditions $\gamma\in(\frac{5}{4},\frac{4}{3}]$. We will employ the paracontrolled method, which was introduced in [GIP15, GST19]. This method is particularly powerful for analyzing the regularity structure of solutions or studying singular SPDEs.
Our strategy as follows. Recall our regularity analysis and difference equation
@@
$=$ $\frac{1}{2}(u^{2})+\xi-\xi$
$=$ $\frac{1}{2}\partial_{x}(v^{2})+\frac{1}{2}\partial_{x}(X^{2})+\partial_{x}(vX).$
Assume that $v$ gains the critical regularity $H^{\frac{1}{2}+}$. By the critical analysis, $\partial_{x}(v^{2})$, $\partial_{x}(X^{2})$ and $\partial_{x}(vX)$ must gain $H^{\frac{1}{2}-\gamma+}$-order regularity at least. But its impossible in the weaker dissipation conditions $\gamma\in(\frac{5}{4},\frac{4}{3}]$ since $\partial_{x}(X^{2})$ and $\partial_{x}(vX)$ can gain $\alpha-1$-order regularity at most. To overcome this difficult, our strategy is to decompose such terms in “ higher-regularity-term" and “ lower- regularity-term" by para-product decomposition (1.9). Correspondingly, we express $v$ as $u^{\sharp}$ and $w$, where $\mathcal{L}u^{\sharp}$ is equal to “ higher-regularity-term" and $\mathcal{L}w$ can counter the “ lower- regularity-term".
We firstly assume that $v$ also belong to $\mathcal{C}^{0+}$. We introduce a method called "para-controlled solution" which developed in [GIP15, GP17] for some part of $u$ notation by $u^{\sharp}$ gaining $H^{\frac{1}{2}+}$ regularity. To get a higher order term in $H^{\frac{1}{2}+}$. Fix distribution $u^{\sharp}$ satisfied
$\mathcal{L}u^{\sharp}=\frac{1}{2}(u^{\sharp})^{2}+R(X,u^{\sharp},u),\quad u^{\sharp}(0)=u_{0}\in W^{\frac{1}{2}+}.$ (4.1)
where $R(X,u^{\sharp},u)$ is to be determined. We assume $R(X,u^{\sharp},u)\in W^{\frac{1}{2}-\gamma+}$. Then by regularity theory, $u^{\sharp}$ gain $W^{\frac{1}{2}+}$. Fix $u=X+w+u^{\sharp}$, then we have
@@
$=$ $\frac{1}{2}(u^{2})+\xi-\xi-\mathcal{L}w$
$=$ $\frac{1}{2}\partial_{x}(w^{2})+\frac{1}{2}\partial_{x}(X^{2})+\partial_{x}(wX)+\partial_{x}(wu^{\sharp})+\partial_{x}(Xu^{\sharp})+(\frac{1}{2}\partial_{x}((u^{\sharp})^{2})-\mathcal{L}w).$
Since $u^{\sharp}$ gains enough regularity, which means $u^{\sharp}\cdot w$ preserve regularity of $w$. The restricted term is $\partial_{x}(w\cdot X)$ and $\partial_{x}(X^{2})$, since they at most gains $\alpha-1$-order regularity which lower than $\frac{1}{2}-\gamma+$ in weak dissipation conditions $\gamma\in(\frac{5}{4},\frac{4}{3}]$. Using para-product decomposition (1.9) and $X\in C_{T}\mathcal{C}^{\alpha}$, we note that $\partial_{x}(X\circ X)$, $\partial_{x}(X\prec w)$ and $\partial_{x}(X\circ w)$ belong to $H^{\frac{1}{2}-\gamma}$, hence the main difficulty is term $\partial_{x}(v\prec X)$ and $\partial_{x}(X\prec X)$.
To solve this, we assume that $v$ satisfied some suitable structure to counter this difficult term. Inspired by para-controlled ansatz, we assume that $w$ as form $u^{\prime}\prec\!\!\!\prec\mathcal{Q}$ and $u^{\mathcal{Q}}:=w+u^{\sharp}$. Assume $u^{\prime}\in\mathcal{W}^{\frac{1}{8}+}(T)$, $\mathcal{Q}\in\mathcal{W}^{\frac{3}{8}+}$ and $u^{\sharp}\in\mathcal{W}^{\frac{1}{2}+}(T)$, then by Lemma 2.5, $w\in\mathcal{W}^{\frac{3}{8}+}(T)$. As $\gamma>\frac{5}{4}$, We have regularity analysis as following
$[\partial_{x}(X\prec X)-X\prec\partial_{x}X]+\frac{1}{2}\partial_{x}(w^{2})+\partial_{x}(wu^{\sharp})+\frac{1}{2}\partial_{x}(X\circ X)\in C_{T}W^{-\frac{3}{4}+},\quad\partial_{x}(X\prec u^{\mathcal{Q}})\in C_{T}W^{-\frac{5}{8}+},$
$\partial_{x}(X\circ u^{\mathcal{Q}})\in C_{T}W^{-\frac{1}{2}+},\quad[u^{\prime}\prec\!\!\!\prec\partial_{x}X-u^{\prime}\prec\partial_{x}X]\in C_{T}W^{-\frac{3}{4}+},\quad u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})\in C_{T}W^{\frac{1}{2}-\gamma+}.$
@@
$R(X,u^{\sharp},u):=$ $\frac{1}{2}\partial_{x}(w^{2})+\partial_{x}(wu^{\sharp})+\partial_{x}(X\circ X)+[\partial_{x}(X\prec X)-X\prec\partial_{x}X]+X\prec\partial_{x}X$
$+\partial_{x}(X\prec u^{\mathcal{Q}}+X\circ u^{\mathcal{Q}}+[\partial_{x}(u^{\mathcal{Q}}\prec X)-u^{\mathcal{Q}}\prec\partial_{x}X]$
$+u^{\mathcal{Q}}\prec\partial_{x}X+u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})+[u^{\prime}\prec\partial_{x}X-u^{\prime}\prec\!\!\!\prec\partial_{x}X]$
Here we see that the purpose of using modified paraproduct operator $f\prec\!\!\!\prec g$ is to ensure that the commutator $\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})-u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}$ aligns with our regularity framework. Fix $u^{\prime}=X+u^{\mathcal{Q}}$, we counter the rough term.
$u^{\prime}\prec\!\!\!\prec\mathcal{L}Q=$ $w\prec\partial_{x}X+X\prec\partial_{x}X$
where we fix $\mathcal{L}\mathcal{Q}=\partial_{x}X$. By equation (4.1), we construct a paracontrolled solution if we can verify $u^{\sharp}$ has enough regularity. In fact, by regularity theory, for $\gamma>\frac{5}{4}$, (4.1) derive the regularity of $u^{\sharp}$ such as
@@ 4.2 Construction of para-controlled solution
We frame this problem as solving a system of equations of $(u^{\prime},u^{\sharp})$. Firstly, we define the equation of $(u^{\prime},u^{\sharp})$ driven by $(\tilde{X},u_{0})$.
$\left\{\begin{aligned} &u^{\prime}=\tilde{X}+u^{\mathcal{Q}}\\ &\mathcal{L}u^{\sharp}=\frac{1}{2}[\partial_{x}(\tilde{X}^{2})-\tilde{X}\prec\partial_{x}\tilde{X}]+\frac{1}{2}\partial_{x}(u^{\mathcal{Q}})^{2}+[\partial_{x}(u^{\mathcal{Q}}\tilde{X})-u^{\mathcal{Q}}\prec\partial_{x}\tilde{X}]+[u^{\prime}\prec\partial_{x}\tilde{X}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})],\end{aligned}\right$. (4.2)
with $u^{\sharp}(0)=u_{0}$.
@@
Let $\alpha\in(\frac{1}{8},\frac{1}{6})$, $\gamma\in(\frac{4}{3},2]$ and $u_{0}\in W^{\frac{1}{2}+\delta}$ for $\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})$. If $\tilde{X}\in\mathcal{W}^{\alpha}(T)$, then there exists a unique solution $(u^{\prime},u^{\sharp})$ of equation (4.2) with positive $T^{*}$ satisfied $(u^{\prime},u^{\sharp})\in\mathcal{W}^{\frac{1}{8}+\delta}(T^{*})\times\mathcal{W}^{\frac{1}{2}+\delta}(T^{*})$.
@@
Firstly, fix $u^{\sharp}(0)=u_{0}$ and $f(u^{\sharp},u^{\prime})=u^{\mathcal{Q}}=u^{\prime}\prec\!\!\!\prec Q+u^{\sharp}$, rewrite (4.2) as following
@@
$\mathcal{L}a=P(t)u_{0}+\frac{1}{2}\partial_{x}(\tilde{X}^{2})-\tilde{X}\prec\partial_{x}\tilde{X}$, $\quad a(0)=0,$
$\mathcal{L}(\mathcal{G}(f(u^{\sharp},u^{\prime})))=[\partial_{x}(u^{\mathcal{Q}}\tilde{X})-u^{\mathcal{Q}}\prec\partial_{x}\tilde{X}$], $\quad\mathcal{G}(f(u^{\sharp},u^{\prime}))(0)=0,$
$\mathcal{L}{\mathcal{B}}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))=\frac{1}{2}\partial_{x}(u^{\mathcal{Q}})^{2},$ $\quad{\mathcal{B}}(f(u^{\sharp},u^{\prime})(0)=0,$
@@
Define $\mathcal{E}_{T}$ as following
@@
Our goal is to find a $T^{*}\leq 1$ such that the conditions of Lemma 2.9 hold in $\mathcal{E}_{T^{*}}$. Firstly, by Lemma 2.5, for $\epsilon=\frac{1}{8}+\delta$ and $\rho=\gamma-\delta$, we have
@@
$\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(T^{*})}.$
@@
$\|f(u^{\sharp},u^{\prime})\|_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}(\|\mathcal{Q}\|_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}+\|\partial_{x}\tilde{X}\|_{\frac{3}{2}-2\gamma+2\delta})$
$\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(T^{*})}.$
fix $T$ small enough, we can make sure that $CT^{1-\frac{\delta}{\gamma}}\|\tilde{X}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T^{*})}<\frac{1}{10}$.
Fix $r=4C\|u_{0}\|_{W^{\frac{1}{2}+\delta}}$, by Lemma A.3 and Lemma B.1, we can choose $T$ small enough such that
@@
$\leq$ $\frac{r}{4}+\frac{r}{2}+\frac{r}{4}\leq r,$
by Lemma B.3 and Lemma B.1, we have similar as above argument, we can choose $T$ small enough such that
$\|g(u^{\prime})\|_{\mathcal{W}^{\frac{1}{2}+\delta}}\leq CT^{\frac{\delta}{\gamma}}\|\mathcal{L}g(u^{\prime})\|_{\mathcal{W}^{\frac{1}{2}-\gamma+2\delta}}\leq C$ $T^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{C_{{T}}\mathcal{C}^{\frac{1}{8}+\delta}}(\|\mathcal{Q}\|_{C_{{T}}W^{\frac{3}{8}+\delta}}+\|\partial_{x}\tilde{X}\|_{C_{T}W^{\frac{3}{8}-\gamma+\delta}})$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{C_{{T}}W^{\alpha}}\leq\frac{1}{16}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}.$
@@
$\|\mathcal{G}(f(u^{\sharp},u^{\prime}))\|_{\mathcal{W}^{\frac{1}{2}+\delta}}\leq$ $CT^{\frac{\delta}{\gamma}}\|\mathcal{L}\mathcal{G}(f(u^{\prime},u^{\sharp}))\|_{\mathcal{W}^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|f(u^{\prime},u^{\sharp})\|_{\mathcal{W}^{\frac{1}{8}+\delta}}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(T)}$
$\leq$ $\frac{1}{16}\|f(u^{\prime},u^{\sharp})\|_{\mathcal{W}^{\frac{1}{8}+\delta}}.$
where we use the fact $\gamma>\frac{3}{2}$. Finally, We can choose $T$ small enough such that
$\|\mathcal{B}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))\|_{\mathcal{W}^{\frac{1}{2}+\delta}}\leq$ $CT^{\frac{\delta}{\gamma}}\|\mathcal{L}\mathcal{B}(f(u^{\prime},u^{\sharp}))\|_{C_{T}W^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|\partial_{x}(f(u^{\prime},u^{\sharp})^{2})\|_{C_{T}W^{\frac{1}{2}+\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|f(u^{\prime},u^{\sharp})^{2}\|_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|f(u^{\prime},u^{\sharp})\|^{2}_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}$
$\leq$ $\frac{1}{8r}(\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}}+c\|y_{n}\|_{\mathcal{W}^{\frac{1}{8}+\delta}})^{2}.$
Fix $T^{*}$ satisfied above all, by Lemma 2.9, there exists $(u^{\sharp},u^{\prime})\in\mathcal{E}_{T^{*}}$ where we prove our proposition. ∎
@@
Choose $\tilde{X}(t)=X(t)$ satisfied (1.5). By the Remark 2.1, we have $\tilde{X}\in\mathcal{W}^{\alpha}({T^{*}})$ for some $\alpha\in(\frac{1}{8},\frac{1}{6})$. Then by Proposition 4.1, there exists $(u^{\prime},u^{\sharp})\in C_{T^{*}}W^{\frac{1}{8}+\delta}\times C_{T^{*}}W^{\frac{1}{2}+\delta}$ satisfied (4.2) for some $\delta>0$. Its not difficult to check that $u=X+u^{\prime}\prec\!\!\!\prec\mathcal{Q}+u^{\sharp}$ is the solution of (1.1) by the analysis in the Section 4.1. Since $u\in C_{T^{*}}W^{\alpha}$, we finish the proof of Theorem 1.3. ∎

### v2

@@ 4.1 Regularity analysis
This section is devoted to isolating a remainder term of Sobolev regularity $\frac{1}{2}+$ under the weaker dissipation condition $\gamma\in(\frac{5}{4},\frac{4}{3}]$. We will achieve this by means of the paracontrolled method, introduced in [GIP15, GST19], a technique especially well‑suited for analyzing the regularity structure of solutions to singular SPDEs.
Our strategy as follows. Recall the difference equation
@@
$=$ $\partial_{x}(u^{2})+\xi-\xi$
$=$ $\partial_{x}(v^{2})+\partial_{x}(X^{2})+2\partial_{x}(vX).$
Assume that $v$ gains the critical regularity $H^{\frac{1}{2}+}$. By the critical analysis, $\partial_{x}(v^{2})$, $\partial_{x}(X^{2})$ and $\partial_{x}(vX)$ must gain $H^{\frac{1}{2}-\gamma+}$-order regularity at least. But it’s impossible in the weaker dissipation conditions $\gamma\in(\frac{5}{4},\frac{4}{3}]$ since $\partial_{x}(X^{2})$ and $\partial_{x}(vX)$ can gain $\alpha-1$-order regularity at most. To overcome this difficult, our strategy is to decompose such terms in “ higher-regularity-term" and “ lower- regularity-term" by para-product decomposition (1.9). Correspondingly, we write
$v=w+u^{\sharp},$ (4.1)
where $\mathcal{L}u^{\sharp}$ “corresponds to the higher-regularity-term" while $\mathcal{L}w$ is designed to absorb the “ lower- regularity-term".
We firstly assume that $v$ also belong to $\mathcal{C}^{0+}$. We introduce a method called "para-controlled solution" which developed in [GIP15, GP17] for some part of $u$ notation by $u^{\sharp}$ gaining $H^{\frac{1}{2}+}$ regularity. To get a higher order term in $H^{\frac{1}{2}+}$. Let distribution $u^{\sharp}$ satisfy equation
$\mathcal{L}u^{\sharp}=\partial_{x}(u^{\sharp})^{2}+R(X,u^{\sharp},u),\quad u^{\sharp}(0)=u_{0}\in W^{\frac{1}{2}+}.$ (4.2)
where $R(X,u^{\sharp},u)$ is to be determined. Assume that $R(X,u^{\sharp},u)\in W^{\frac{1}{2}-\gamma+}$. Then by regularity theory, $u^{\sharp}$ gain $W^{\frac{1}{2}+}$. Fix $u=X+w+u^{\sharp}$, then we have
@@
$=$ $\partial_{x}(u^{2})+\xi-\xi-\mathcal{L}w$
$=$ $\partial_{x}(w^{2})+\partial_{x}(X^{2})+2\partial_{x}(wX)+2\partial_{x}(wu^{\sharp})+2\partial_{x}(Xu^{\sharp})+(\partial_{x}((u^{\sharp})^{2})-\mathcal{L}w).$
Since $u^{\sharp}$ possesses sufficient regularity, the product $u^{\sharp}\cdot w$ preserve regularity of $w$. The problematic terms are $\partial_{x}(w\cdot X)$ and $\partial_{x}(X^{2})$, as they can gain at most $\alpha-1$-order regularity which below the required $\frac{1}{2}-\gamma+$ in weak dissipation conditions $\gamma\in(\frac{5}{4},\frac{4}{3}]$. Using para-product decomposition (1.9) and the fact that $X\in C_{T}\mathcal{C}^{\alpha}$, we observe that $\partial_{x}(X\circ X)$, $\partial_{x}(X\prec w)$ and $\partial_{x}(X\circ w)$ all belong to $H^{\frac{1}{2}-\gamma}$. The main difficulty therefore lies in the terms $\partial_{x}(v\prec X)$ and $\partial_{x}(X\prec X)$.
To handle this, we assume that $v$ possesses a suitable structure that can absorb the problematic terms. Motivated by the paracontrolled ansatz, we write
$w=u^{\prime}\prec\!\!\!\prec\mathcal{Q},\quad u^{\mathcal{Q}}:=w+u^{\sharp}.$ (4.3)
with the assumption $u^{\prime}\in\mathcal{W}^{\frac{1}{8}+}(T)$, $\mathcal{Q}\in\mathcal{W}^{\frac{3}{8}+}$ and $u^{\sharp}\in\mathcal{W}^{\frac{1}{2}+}(T)$. By Lemma 2.5, this implies $w\in\mathcal{W}^{\frac{3}{8}+}(T)$. For $\gamma>\frac{5}{4}$, the regularity analysis proceeds as follows
$2[\partial_{x}(X\prec X)-X\prec\partial_{x}X]+\partial_{x}(w^{2})+2\partial_{x}(wu^{\sharp})+\partial_{x}(X\circ X)\in C_{T}W^{-\frac{3}{4}+},\quad 2\partial_{x}(X\prec u^{\mathcal{Q}})\in C_{T}W^{-\frac{5}{8}+},$
$2\partial_{x}(X\circ u^{\mathcal{Q}})\in C_{T}W^{-\frac{1}{2}+},\quad 2[u^{\prime}\prec\!\!\!\prec\partial_{x}X-u^{\prime}\prec\partial_{x}X]\in C_{T}W^{-\frac{3}{4}+},\quad u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})\in C_{T}W^{\frac{1}{2}-\gamma+}.$
@@
$R(X,u^{\sharp},u):=$ $\partial_{x}(w^{2})+2\partial_{x}(wu^{\sharp})+\partial_{x}(X\circ X)+2[\partial_{x}(X\prec X)-X\prec\partial_{x}X]$
$+2\partial_{x}(X\prec u^{\mathcal{Q}}+X\circ u^{\mathcal{Q}}+[\partial_{x}(u^{\mathcal{Q}}\prec X)-u^{\mathcal{Q}}\prec\partial_{x}X]$
$+u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})+[u^{\prime}\prec\partial_{x}X-u^{\prime}\prec\!\!\!\prec\partial_{x}X].$
Here we see that the purpose of using modified paraproduct operator $f\prec\!\!\!\prec g$ is to ensure that the commutator $\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})-u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}$ aligns with our regularity framework. Let $u^{\prime}=X+u^{\mathcal{Q}}$, we counter the rough term.
$u^{\prime}\prec\mathcal{L}Q=$ $2X\prec\partial_{x}X+2u^{\mathcal{Q}}\prec\partial_{x}X,$
where the last term has higher regularity. By equation (4.2), we construct a paracontrolled solution if we can verify $u^{\sharp}$ has enough regularity. In fact, by regularity theory, for $\gamma>\frac{5}{4}$, (4.2) derive the regularity of $u^{\sharp}$ such as
@@ 4.2 Construction of para-controlled solution
We frame this problem as solving a system of equations of $(u^{\prime},u^{\sharp})$. Given an equation of $(u^{\prime},u^{\sharp})$ driven by $(\tilde{X},u_{0})$.
$\left\{\begin{aligned} &u^{\prime}=2\tilde{X}+2u^{\mathcal{Q}}\\ &\mathcal{L}u^{\sharp}=[\partial_{x}(\tilde{X}^{2})-\tilde{X}\prec\partial_{x}\tilde{X}]+\partial_{x}(u^{\mathcal{Q}})^{2}+2[\partial_{x}(u^{\mathcal{Q}}\tilde{X})-u^{\mathcal{Q}}\prec\partial_{x}\tilde{X}]+2[u^{\prime}\prec\partial_{x}\tilde{X}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})],\end{aligned}\right$. (4.4)
with $u^{\sharp}(0)=u_{0}$. We have the following proposition
@@
Let $\alpha\in(\frac{1}{8},\frac{1}{6})$, $\gamma\in(\frac{5}{4},\frac{4}{3}]$ and $u_{0}\in W^{\frac{1}{2}+\delta}$ for $0<\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})$. If $\tilde{X}\in\mathcal{W}^{\alpha}(\bar{T})$ for some $\bar{T}>0$, then there exists a unique solution $(u^{\prime},u^{\sharp})$ of equation (4.4) with positive $T^{*}$ satisfied $(u^{\prime},u^{\sharp})\in\mathcal{W}^{\frac{1}{8}+\delta}(T^{*})\times\mathcal{W}^{\frac{1}{2}+\delta}(T^{*})$.
@@
Firstly, let $u^{\sharp}(0)=u_{0}$ and $f(u^{\sharp},u^{\prime})=u^{\mathcal{Q}}=u^{\prime}\prec\!\!\!\prec Q+u^{\sharp}$, we can rewrite (4.4) in mild form
@@
$\mathcal{L}a=P(t)u_{0}+[\partial_{x}(\tilde{X}^{2})-2\tilde{X}\prec\partial_{x}\tilde{X}$], $\quad a(0)=0,$
$\mathcal{L}(\mathcal{G}(f(u^{\sharp},u^{\prime})))=2[\partial_{x}(u^{\mathcal{Q}}\tilde{X})-u^{\mathcal{Q}}\prec\partial_{x}\tilde{X}$], $\quad\mathcal{G}(f(u^{\sharp},u^{\prime}))(0)=0,$
$\mathcal{L}{\mathcal{B}}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))=\partial_{x}(u^{\mathcal{Q}})^{2},$ $\quad{\mathcal{B}}(f(u^{\sharp},u^{\prime})(0)=0,$
@@
Define $\mathcal{E}_{T}$ as follows
@@
Our goal is to find a $T^{*}\leq 1\wedge\bar{T}$ such that the conditions of Lemma 2.9 hold in $\mathcal{E}_{T^{*}}$. Firstly, by Lemma 2.5, for $\epsilon=\frac{1}{8}+\delta$, $\rho=\gamma-\delta$ and $\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})$, we have
@@
$\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(\bar{T})}.$
@@
$\|f(u^{\sharp},u^{\prime})\|_{\mathcal{W}^{\frac{3}{2}-\gamma+\delta}(T)}\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}(\|\mathcal{Q}\|_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}+\|\partial_{x}\tilde{X}\|_{\frac{3}{2}-2\gamma+2\delta})$
$\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(\bar{T})}.$
Fixing $T$ small enough, we can make sure that $CT^{1-\frac{\delta}{\gamma}}\|\tilde{X}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T^{*})}<\frac{1}{30}$.
For fixed $r=16(1\vee C)\max(\|u_{0}\|_{W^{\frac{1}{2}+\delta}},\|\tilde{X}\|_{\mathcal{W}^{\alpha}(\bar{T})})$, by Lemma A.2 and Lemma B.1, we can choose $T\leq 1\wedge\bar{T}$ small enough such that
@@
$\leq$ $\frac{r}{16}+\frac{r}{8}+\frac{r}{16}\leq r.$
By Lemma B.3 and Lemma B.1, we have similar as above argument, we can choose $T\leq 1\wedge\bar{T}$ small enough such that
$\|g(u^{\prime})\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}\leq CT^{\frac{\delta}{\gamma}}\|\mathcal{L}g(u^{\prime})\|_{C_{T}W^{\frac{1}{2}-\gamma+2\delta}}\leq C$ $T^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}(\|\mathcal{Q}\|_{C_{{T}}W^{\frac{3}{8}+\delta}}+\|\partial_{x}\tilde{X}\|_{C_{T}W^{\frac{3}{8}-\gamma+\delta}})$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}\leq\frac{1}{16}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}.$
@@
$\|\mathcal{G}(f(u^{\sharp},u^{\prime}))\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}\leq$ $CT^{\frac{\delta}{\gamma}}\|\mathcal{L}\mathcal{G}(f(u^{\sharp},u^{\prime}))\|_{C_{T}W^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|f(u^{\sharp},u^{\prime})\|_{\mathcal{W}^{\frac{1}{8}+\delta}}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(\bar{T})}$
$\leq$ $\frac{1}{16}\|f(u^{\sharp},u^{\prime})\|_{\mathcal{W}^{\frac{1}{8}+\delta}}.$
where we use the fact $\gamma>\frac{3}{2}$. Finally, let $X=\mathcal{W}^{\frac{3}{2}-\gamma+\delta}(T)$, we have $\mathcal{W}^{\frac{1}{2}+\delta}(T)\subset\mathcal{W}^{\frac{3}{2}-\gamma+\delta}(T)\subset\mathcal{W}^{\frac{1}{8}+\delta}(T)$, Choosing $T$ small enough such that
$\|\mathcal{B}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}\leq$ $CT^{\frac{\delta}{\gamma}}\|\mathcal{L}\mathcal{B}(f(u^{\sharp},u^{\prime}))\|_{C_{T}W^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|\partial_{x}(f(u^{\sharp},u^{\prime})^{2})\|_{C_{T}W^{\frac{1}{2}-\gamma+\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|f(u^{\sharp},u^{\prime})\|^{2}_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}$
$\leq$ $\frac{1}{16}\|f(u^{\sharp},u^{\prime})\|^{2}_{\mathcal{W}^{\frac{3}{2}-\gamma+\delta}(T)}.$
Setting $T^{*}$ to satisfy all the conditions above, by Lemma 2.9, there exists $(u^{\sharp},u^{\prime})\in\mathcal{E}_{T^{*}}$ where we prove our proposition. ∎
@@
Let $\tilde{X}(t)=X(t)$ where $X(t)$ satisfies equation (1.5), by the Remark 2.1, we have $\tilde{X}\in\mathcal{W}^{\alpha}({\bar{T}})$ for any fixed $\bar{T}>0$ and $\alpha\in(\frac{1}{8},\frac{\gamma}{2}-\frac{1}{2})$. Then by Proposition 4.1, there exists $(u^{\prime},u^{\sharp})\in C_{T^{*}}W^{\frac{1}{8}+\delta}\times C_{T^{*}}W^{\frac{1}{2}+\delta}$ satisfied (4.4) for $\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})<\frac{\gamma}{2}-\frac{5}{8}$. It’s not difficult to check that $u=X+u^{\prime}\prec\!\!\!\prec\mathcal{Q}+u^{\sharp}$ is the solution of (1.1) by the analysis in the Section 4.1. Since $u\in C_{T^{*}}W^{\alpha}$, we finish the proof of Theorem 1.3. ∎